# Boosting Mixed-Initiative Co-Creativity in Game Design: A Tutorial

### v1

@@ 4.2.9. Germinate
Task assignment: Both agents are focused on creating a game that expresses specific rhetorical goals. However, they cannot engage in the same type of actions in the co-creative process. While the computational agent is capable of generating complete games, the human agent can only define cards in a way that, hopefully, their intent is understood and conveyed in the games based solely on these guidelines. The contributions of the computational agent do not directly affect the subsequent actions of the human agent.
@@
Intervention pace: Each agent has their own defined shift: the human agent is responsible for setting the cards, while the computational agent generates batches of games. These shifts tend to be lengthy. The system takes a considerable amount of time to generate the first batch of games, leaving the user waiting without any feedback during this period. Due to the extended duration of the computational agent’s turn, it is likely that the human agent’s turn will also be long, as they may choose to make multiple modifications at once. Moreover, as the content being created transitions from properties described in cards to full games generated by the system, the content leap is significant.
@@
Explainability: This approach lacks extensive explainability, though it offers more than the previous works. In addition to using cards to express their intentions, the human co-creator can provide further clarification by adding tags to entity or resource cards. For instance, they may include a “Stress” resource with the tag “player thinks it’s bad”, enabling the computational agent to better grasp their vision. Conversely, each game generated by the computational agent includes its rules displayed alongside in the form of cards, similar to the ones used by the human agent to define their intent. However, the presentation of these cards alone is often insufficient, and the human agent still struggles to comprehend the exact rules of the game and the message it aims to convey. A significant issue with this tool lies in its lack of explainability when the system detects inconsistencies in the user’s intent and is unable to generate games. In such cases, the system merely states that it cannot generate the games, without explaining the existing contradictions. This hinders the co-creative process, leaving the user unaware of what changes are necessary to resolve the issue.
@@ 4.2.10. Overview
As previously mentioned, this is not an exhaustive analysis, and different individuals may arrive at distinct conclusions regarding the level of MI-CCy in each tool. Nevertheless, it enables us to compare the approaches and identify relevant features in some that may be lacking in others. Additionally, it enables us to pinpoint characteristics common to most approaches, which can be addressed when developing future MI-CCy approaches in game design. In Figure 11, we provide an overview classification of each tool based on the MI-CCy Quantifier criteria. We can observe that the works that stand out with a higher level of MI-CCy are the Sentient Sketchbook and 1001 Nights. By examining Figure 11 alone, several patterns become discernible. Specifically, the initial setting (i.e., defining the problem at hand, the conceptual space, and the starting point) tends to be skewed towards one of the agents. The evaluation of the process, artifacts, and creativity of the agents is primarily conducted by the human co-creator. Additionally, the final decision is always up to the human agent. Lastly, there is a general lack of explainability, both from the computational agent to the human agent and vice versa.
As mentioned earlier, game design is a broad discipline, so these approaches can address several types of game content at different complexities. Furthermore, there are a variety of techniques that can be used depending on the goals and intended functionality. We can identify some trends in the techniques employed in these MI-CCy approaches. Evolutionary algorithms are among the most commonly used techniques in this context. For instance, the Sentient Sketchbook utilizes genetic algorithms, while both the AI Spaceship Generator and Story Designer employ MAP-Elites. The use of machine learning algorithms is also prevalent, as seen in the RL Brush, which relies on reinforcement-learning-based models, and the Sprite Editor for Pixel Art Characters, which uses generative adversarial networks. 1001 Nights makes use of Natural Language Processing, a technique often seen in co-creative writing tools. Among the tools we have analyzed, there are also some less conventional content generation techniques. That is the case for the miWFC, which applies the Wave Function Collapse algorithm, and Germinate, which makes use of answer set programming to generate games. The Pitako tool also stands out for its recommendation-based approach through frequent itemset mining.
The survey conducted by Lai et al. provides a more extensive review of the techniques used in mixed-initiative systems for game content creation. From our analysis, we contend that the underlying algorithms in these systems do not have as significant an impact on the level of MI-CCy as the means through which agents can contribute with creative input and communicate with one another. Nonetheless, we observe that techniques that encourage the computational agent to generate more novel and interesting content enable it to play a more significant role in the process.
In addition to the algorithms and techniques used, it is crucial to overview how agents interact within the creative process. Regarding the computational agent, it is frequent for their contributions to involve presenting a set of suggestions from which the human agent can choose to accept one. This trend is observed in six out of the nine systems assessed using the MI-CCy Quantifier. Only in two works, namely miWFC and 1001 Nights, does the computational agent directly add content to the main creative artifact. Finally, in the Sprite Editor for Pixel Art Characters, the computational agent offers a single suggestion at a time for the user to accept or reject. Conversely, the human agent can add content directly to the final artifact in seven out of the nine frameworks, underscoring a clear distinction from their computational partner. The only exceptions are the AI Spaceship Generator, where the human agent is limited to selecting suggestions, and Germinate, where they can define cards instead of directly designing the game.
@@
Figure 11. Overview of the analysis of the degree of mixed-mixed-initiative co-creativity (MI-CCy) for each work using the MI-CCy Quantifier.
@@ 5. Discussion and Open Issues
The works presented in Section 3 and further explored in Section 4.2 express a fair sample of the current progress in the development of mixed-initiative approaches to game design. Several other works were analyzed but not elected as use cases for this tutorial. Despite the diversity of recent works, there are some open issues. Below we describe the main ones identified:
@@
Most approaches still fail to promote seamless collaboration between the human and the computer. In most cases, the human and the computational agent are still performing separated steps at a time and, therefore, not learning much from each other. For instance, in Germinate (Kreminski et al., 2020), the user sets the initial properties, and then the computer generates games according to those defined properties. Subsequently, the user can analyze the generated content and make changes to the properties, followed by the computer generating more games. Various steps of modifying properties according to the new knowledge acquired from the output can result in a reasonable co-creative process. However, the rigid alternating of tasks remains noticeable, and the agents have a hard time engaging in a productive exchange of ideas that allows for the ideas to be smoothly blended.
@@
There is an unbalanced amount of initiative and creative input between the human and the computational agent. It is observable that some works favor computational initiative, while others favor human initiative. In the work of Gallotta et al. (Gallotta et al., 2023) the computational agent generates the first population of spaceships and has almost full control over what is presented. The human counterpart has low initiative and creative input as they can only select spaceships throughout the process. In contrast, in the Sprite Editor for Pixel Art Characters (dos Santos Coutinho and Chaimowicz, 2023) the computer is not as creative as the human, as the human partner creates the base character design, and the system always tries to match their vision during the course of co-creation. Compounding this issue, there are generally few instances where the computational agent is completely spontaneous, i.e., cases where it can react without requiring any trigger from the human counterpart. With the exception of when the computer presents an initial setting, if the human agent does not take any action, the computer remains inactive as well.
@@
There is a lack of research in some domains of game design. Level design is the type of game content most tackled by current approaches. Although relevant, the high focus on map creation results in other critical aspects of game design receiving less consideration. Game content such as characters, sound, user interface, or complex interactions between game elements tend to be overlooked. Narrative design and the creation of visual assets, such as 2D or 3D models, although prevalent in many co-creative tools across broader domains, do not receive as much attention within the specific realm of video games.
@@
We recommend that individuals interested in developing novel mixed-initiative co-creative tools for game design should assess whether their system addresses any of these gaps. The MI-CCy Quantifier can serve as a valuable tool for evaluating how well your developing system performs against each criterion and what improvements it brings to the current state-of-the-art. We note that the MI-CCy Quantifier is not a closed framework, meaning that additional dimensions can potentially be incorporated beyond the ones we have discussed here. Usability is one such example. A user-friendly interface is crucial for facilitating interaction between the human and computational agents. While testing these tools, we observed variations in usability, with some demonstrating superior usability compared to others. This directly impacted our willingness to engage with the system and the quality of the co-creative process. A tool with poor usability generally results in a lower level of co-creativity.
@@ 6. Conclusion
In this paper, we provided guidance to researchers, developers, and game designers interested in implementing new approaches to game design focused on a high level of mixed-initiative co-creativity. To do so, we presented existing works on MI-CCy in game design targeted at several classes of game content. We introduced the MI-CCy Quantifier, a directive framework for analyzing works based on their level of MI-CCy in a descriptive and visually quantifiable manner. We illustrated its usage by applying it to nine of the presented frameworks. Through this analysis, we identified patterns in current approaches and characteristics that favor some over others. Furthermore, we highlight prevalent open issues in the state-of-the-art, which we propose as key points to address when conceiving new MI-CCy frameworks for game design. In particular, we recommend that people working on forthcoming approaches focus on promoting a seamless and balanced partnership between the human agent and the computational agent. We also emphasize the need to address overlooked types of game content, to aim for greater versatility, and to enhance explainability. Finally, it is crucial to establish methods for evaluating MI-CCy approaches.
@@ Acknowledgements.
This work is funded by the FCT – Foundation for Science and Technology, I.P./MCTES through national funds (PIDDAC), within the scope of CISUC R&D Unit – UIDB/00326/2020 or project code UIDP/00326/2020. The first author is funded by Foundation for Science and Technology (FCT), Portugal, under the grant 2021.07047.BD.

### v2

@@ 4.2.5. Task assignment
Germinate (G): Both agents are focused on creating a game that expresses specific rhetorical goals. However, they cannot engage in the same type of actions in the co-creative process. While the computational agent is capable of generating complete games, the human agent can only define cards in a way that, hopefully, their intent is understood and conveyed in the games based solely on these guidelines. The contributions of the computational agent do not directly affect the subsequent actions of the human agent.
Figure 16. Overview of the ranking of the Task assignment for each work.
4.2.6. Intervention pace
AI Spaceship Generator (SG): The agents have clearly defined turns. Even if the user is quick to select one (or none) spaceship to evolve from in the next iteration, there is a long waiting period where the user cannot do anything nor receive feedback until the system completes the new generation. Furthermore, the system responds with new complete spaceship designs, as opposed to short modifications to the selected spaceship or more open-ended prototypes.
Sprite Editor for Pixel Art Characters (SE): The participants involved in the co-creative process can engage in a fast-paced interaction loop. Despite having well-defined turns for each participant (with the system only intervening upon user request), the response is immediate. In most cases, the suggestions from the system do not represent an extensive leap in content and are closely aligned with the user-generated content. Since the human agent has more control in the process, they can generate from none to a substantial amount of new content between the interventions of the computational agent. Although the turns are short, the system cannot intervene spontaneously.
Sentient Sketchbook (SS): The co-creators engage in a fast intervention pace, with short turns involving the addition of small portions of content over time. In this case, the human agent edits a single tile (or a small area) at a time, and the computational agent intervenes with each change made. The content leap between the main map sketch and the computer’s suggestions is also not too substantial. There is only a larger gap observed at the beginning of the co-creative process as the system generates suggestions with sufficient content to be playable. In essence, the agents provide immediate and spontaneous feedback.
RL Brush (RB): The intervention pace of the computational agent is highly dependent on the values of the parameters “step” and “tool radius”. When the “step” value is low and the “tool radius” is small, the feedback of the system is nearly instantaneous. However, if these values are increased, the system’s response time also increases. Nevertheless, we consider this waiting time to be reasonable. Both the human and the computer edit one or a few tiles at a time, resulting in small content leaps. Consequently, the agents are not overwhelmed by a large amount of new information all at once, allowing them to gradually build the level and merge their ideas more easily.
miWFC (mW): The co-creators have well-defined shifts. When the human agent is editing the image, the computational agent waits and only intervenes upon the human partner’s request. Despite this, the computational agent can generate content immediately, enabling rapid interaction between the agents. On the other hand, for the human designer to manually edit the canvas, they must switch screens, inevitably slowing down the intervention pace. The amount of content produced by each agent during their turn varies according to how the human agent chooses to interact with the computational agent. The computational agent can generate anything from a single tile to a complete bitmap, depending on the extent to which the human partner allows.
Pitako (P): The pace and spontaneity of co-creators’ interventions depend on whether they are working on creating sprites or interactions. In the case of sprites, the turns are short, with the computational agent typically providing suggestions after each user intervention. Both parties focus on adding one sprite at a time. However, creating a sprite involves not only defining its graphical representation but also its behavior, which may require specifying several properties. Thus, we consider the content leap to be moderate. In the case of interactions, the system does not respond spontaneously to the contributions of the human creator, providing recommendations only upon request and solely based on the sprite set. Consequently, the user receives no feedback as they add interactions to the game.
Story Designer (SD): The system is not swift enough to generate suggestions in real-time as the user makes alterations, often resulting in delayed responses. Moreover, to view the suggestions generated by the computational agent, the human partner needs to inspect them one by one, thereby slowing down the intervention pace. As for the extent of the content leap, it varies from suggestion to suggestion. For instance, suggestions with a higher value in the “step” dimension will entail a more significant content leap.
1001 Nights (1N): In this approach, the agents have precise turns where each one contributes a portion of the story alternately. Although the turns are relatively short, we consider the content leap to be moderate, especially on the side of the computational agent, as it can generate more content in its turn. While the human agent may write one or two sentences, the computational agent can write an entire paragraph uninterrupted. The intervention pace would be faster if the agents could react while their partner is writing.
@@ 4.2.6. Intervention pace
Germinate (G): Each agent has their own defined shift: the human agent is responsible for setting the cards, while the computational agent generates batches of games. These shifts tend to be lengthy. The system takes a considerable amount of time to generate the first batch of games, leaving the user waiting without any feedback during this period. Due to the extended duration of the computational agent’s turn, it is likely that the human agent’s turn will also be long, as they may choose to make multiple modifications at once. Moreover, as the content being created transitions from properties described in cards to full games generated by the system, the content leap is significant.
Figure 17. Overview of the ranking of the Intervention pace for each work.
4.2.7. Explainability
AI Spaceship Generator (SG), Sprite Editor for Pixel Art Characters (SE), RL Brush (RB): There is no explainability. Neither the human nor the computer has the means to express their intentions and justify their contributions.
Sentient Sketchbook (SS): The explainability is severely limited. The computational agent makes a feeble attempt to justify why its suggestions may be preferable by displaying their respective fitness scores compared to the main map sketch. Thus, in the case of suggestions aimed at optimizing fitness scores, the human agent has a means to understand some of the improvements the suggestions offer to gameplay. However, these properties solely describe functional characteristics and provide no insight into the creativity of the map design. Conversely, the human designer lacks any means to explain their intentions and preferences.
miWFC (mW): There is little explainability between the human and the computational agent. All the user knows is that the system attempts to create content similar to the template used as an example, and according to the tile properties defined by the user. Additionally, the system’s display of an animation of the generative process that the user can analyze step by step allows for a better understanding of the creative process of the computational agent. Nevertheless, the human agent has no other means of gaining further insight into the computational agent’s decisions, nor does the computational agent have a way to comprehend the creative goals and preferences of the human agent.
Pitako (P): Explainability is almost non-existent. The computational agent provides a small hint regarding the novelty of its recommendations by displaying the confidence level, leaving it to the discretion of the human partner whether they want to create a game that closely resembles existing ones or opt for a set of sprites and interactions that is more distinct. However, it does not provide any further information about its creative process, and vice versa.
Story Designer (SD): Similar to the Sentient Sketchbook, the system displays the fitness, interestingness, and coherence scores of suggestions in relation to the main narrative graph. This provides some insight into the specific optimizations certain suggestions offer and why they might be preferable. However, this level of explainability remains insufficient, and there is no means for the human agent to reciprocally provide justification.
1001 Nights (1N): Since the game’s objective is for the player to prompt the King to mention specific weapons, the player may sometimes become puzzled when the system’s response lacks any of the keywords. In such cases, the system provides the player with a hint that they should tell a story more centered around swords and magic. However, this explanation is identical for every instance in which the King fails to mention any of the items. It also provides no further insight into the overall creative process of the AI in story writing. Consequently, explainability is almost absent.
@@ 4.2.7. Explainability
Germinate (G): This approach lacks extensive explainability, though it offers more than the previous works. In addition to using cards to express their intentions, the human co-creator can provide further clarification by adding tags to entity or resource cards. For instance, they may include a “Stress” resource with the tag “player thinks it’s bad”, enabling the computational agent to better grasp their vision. Conversely, each game generated by the system includes its rules displayed alongside in the form of cards. However, the presentation of these cards alone is often insufficient, and the human agent still struggles to comprehend the exact rules of the game and the message it aims to convey. A significant issue with this tool lies in its lack of explainability when it detects inconsistencies in the user’s intent. In such cases, the system merely states that it cannot generate the games, without explaining the existing contradictions. This hinders the co-creative process, leaving the user unaware of what changes are necessary to resolve the issue.
Figure 18. Overview of the ranking of the Explainability for each work.
@@ 4.2.8. Overview
As previously mentioned, this is not an exhaustive analysis, and different individuals may arrive at distinct conclusions regarding the level of MI-CCy in each tool. Nevertheless, it enables us to compare the approaches and identify relevant features in some that may be lacking in others. Additionally, it enables us to pinpoint characteristics common to most approaches, which can be addressed when developing future MI-CCy approaches in game design. In Figure 19, we provide an overview classification of each tool based on the MI-CCy Quantifier criteria. We can observe that the works that stand out with a higher level of MI-CCy are the Sentient Sketchbook and 1001 Nights. By examining Figures 12 to 18 or Figure 19 alone, several patterns become discernible. Specifically, the initial setting tends to be skewed towards one of the agents. The evaluation of the process, artifacts, and creativity of the agents is primarily conducted by the human co-creator. Additionally, the final decision is always up to the human agent. Lastly, there is a general lack of explainability, both from the computational agent to the human agent and vice versa.
@@
Figure 19. Overview of the analysis of the degree of mixed-mixed-initiative co-creativity for each work using the MI-CCy Quantifier.
As mentioned earlier, game design is a broad discipline, so these approaches can address several types of game content with different complexities. Furthermore, there are a variety of techniques that can be used depending on the goals and intended functionality. We can identify trends in the techniques employed in these MI-CCy tools, such as the use of evolutionary algorithms. For instance, the Sentient Sketchbook uses genetic algorithms, while both the AI Spaceship Generator and Story Designer employ MAP-Elites. Machine learning algorithms are also prevalent, as seen in the RL Brush, which relies on reinforcement-learning-based models, and the Sprite Editor for Pixel Art Characters, which uses Generative Adversarial Networks. 1001 Nights makes use of Natural Language Processing, often seen in co-creative writing tools. Among the tools we have analyzed, there are also some less conventional content generation techniques. Namely, the miWFC applies the Wave Function Collapse algorithm, Germinate makes use of answer set programming to generate games, and Pitako applies a recommendation-based approach through frequent itemset mining. The survey conducted by Lai et al. (Lai et al., 2022) provides a more extensive review of the techniques used in mixed-initiative systems for game content creation. From our analysis, we contend that the underlying algorithms in these systems do not have as significant an impact on the level of MI-CCy as the means through which agents can contribute with creative input and communicate with one another. Nonetheless, we observe that techniques that encourage the computer to generate more novel and interesting content enable it to play a more significant role in the process.
In addition to the algorithms and techniques used, it is crucial to review how agents interact within the creative process. Regarding the computational agent, it is frequent for its contributions to involve presenting a set of suggestions from which the human agent can choose to accept one. This trend is observed in six out of the nine systems assessed using the MI-CCy Quantifier. Only in two works, namely miWFC and 1001 Nights, does the computational agent directly add content to the main creative artifact. Finally, in the Sprite Editor for Pixel Art Characters, the computational agent offers a single suggestion at a time for the user to accept or reject. Conversely, the human agent can add content directly to the final artifact in seven out of the nine frameworks, underscoring a clear distinction from their computational partner. The only exceptions are the AI Spaceship Generator, where the human agent is limited to selecting suggestions, and Germinate, where they can define cards instead of directly designing the game.
@@ 5. Discussion and Open Issues
The works presented in Section 3 and further explored in Section 4.2 express a fair sample of the current progress in the development of mixed-initiative approaches to game design. Several other works were analyzed but not selected as use cases for this tutorial. Despite the diversity of recent works, there are some open issues. Below we describe the main ones identified:
@@
Most approaches still fail to promote seamless collaboration between the human and the computer. In most cases, the human and the computational agent are still performing separate steps at a time and, therefore, not learning much from each other. For instance, in Germinate (Kreminski et al., 2020), the user sets the initial properties, and then the computer generates games according to those defined properties. Subsequently, the user can analyze the generated content and make changes to the properties, followed by the computer generating more games. Various steps of modifying properties according to the new knowledge acquired from the output can result in a reasonable co-creative process. However, the rigid alternation of tasks remains noticeable, and the agents have a hard time engaging in a productive exchange of ideas that allows for the ideas to be smoothly blended.
@@
There is an unbalanced amount of initiative and creative input between the human and the computational agent. It is observable that some works favor computational initiative, while others favor human initiative. In the work of Gallotta et al. (Gallotta et al., 2023), the computational agent generates the first population of spaceships and has almost full control over what is presented. The human counterpart has low initiative and creative input, as they can only select spaceships throughout the process. In contrast, in the Sprite Editor for Pixel Art Characters (Coutinho and Chaimowicz, 2024), the computer is not as creative as the human, as the human partner creates the base character design, and the system always tries to match their vision during the course of co-creation. Compounding this issue, there are generally few instances where the computational agent is completely spontaneous, i.e., cases where it can react without requiring any trigger from the human counterpart. With the exception of when the computer presents an initial setting, if the human agent does not take any action, the computer remains inactive as well.
@@
There is a lack of research in some domains of game design. Level design is the type of game content most tackled by current approaches. Although relevant, the high focus on map creation results in other critical aspects of game design receiving less consideration. Game content such as characters, sound, user interface, or complex interactions between game elements tends to be overlooked. Narrative design and the creation of visual assets, such as 2D or 3D models, although prevalent in many co-creative tools across broader domains, do not receive as much attention within the specific realm of video games.
@@
We recommend that individuals interested in developing novel mixed-initiative co-creative tools for game design should assess whether their system addresses any of these gaps. The MI-CCy Quantifier can serve as a valuable tool by providing a structured way to evaluate how well your developing system performs against each criterion and what improvements it brings to the current state of the art. Its utility is leveraged when the framework is used to analyze multiple co-creative systems. The comparison of how the systems differ across the MI-CCy Quantifier parameters allows us to identify trends and outliers, offering a measurable outlook on the challenges faced by current approaches. Another point to highlight is that the analysis is always subjective to the individual experience of the user conducting it. If the MI-CCy Quantifier is used as a tool for individual reflection, this is not an issue. However, if the goal is to draw more objective conclusions about a system’s classification, a study should be conducted with multiple users to identify trends in their assessments.
We note that the MI-CCy Quantifier is not a closed framework, meaning that additional dimensions can potentially be incorporated beyond the ones we have discussed here. Usability is one such example. A user-friendly interface is crucial for facilitating interaction between the human and computational agents. While testing these tools, we observed variations in usability, with some demonstrating superior usability compared to others. This directly impacted our willingness to engage with the system and the quality of the co-creative process. A tool with poor usability generally results in a lower level of co-creativity.
On a final note, while this tutorial primarily focuses on providing guidelines for developing and assessing mixed-initiative co-creativity in the context of game design tools, we recognize that there are critical ethical issues surrounding AI usage that deserve attention. Some key concerns include how AI might affect intellectual property, especially as the division between human and computational-generated content becomes blurred, raising questions about authorship and ownership. There are also possible impacts on labor dynamics, as AI tools can change the roles and responsibilities of human creators, potentially affecting job security, fair compensation, and how human creativity is valued. On a broader level, AI systems might reinforce existing biases or reduce chances for human collaboration. These are complex issues that deserve thoughtful consideration, especially as AI becomes more involved in creative fields. Although this paper does not directly address these ethical challenges, we consider them important areas for future research to ensure that MI-CCy develops in a fair and responsible way.
@@ 6. Conclusion
In this paper, we provided guidance to researchers, developers, and game designers interested in implementing new approaches to game design focused on a high level of mixed-initiative co-creativity. To do so, we presented existing works on MI-CCy in game design targeted at several classes of game content. We introduced the MI-CCy Quantifier, a directive framework for analyzing works based on their level of MI-CCy in a descriptive and visually gradable manner. We illustrated its usage by applying it to nine of the presented frameworks. Through this analysis, we identified patterns in current approaches and characteristics that favor some over others. Furthermore, we highlight prevalent open issues in the state of the art, which we propose as key points to address when conceiving new MI-CCy frameworks for game design. In particular, we recommend that people working on forthcoming approaches focus on promoting a seamless and balanced partnership between the human agent and the computational agent. We also emphasize the need to address overlooked types of game content, to aim for greater versatility, and to enhance explainability. Finally, it is crucial to establish methods for evaluating MI-CCy approaches.
@@ Acknowledgements.
This work is partially financed through national funds by FCT – Fundação para a Ciência e a Tecnologia, I.P., in the framework of the Project UIDB/00326/2025 and UIDP/00326/2025; and also supported by the Portuguese Recovery and Resilience Plan (PRR) through project C645008882-00000055, Center for Responsible AI. The first author is funded by Fundação para a Ciência e a Tecnologia (FCT), Portugal, under the grant 2021.07047.BD.